A ball of mass 1.0 kg moving east with a speed of 2.0 m/s collideshead-on with a 2.0-kg ball at rest. If the collision is perfectly inelastic,what will be the speed and direction of each ball after the collision? (Show all Work)
Answer #1

Using conservation of momentum

m1 v1 = ( m1 + m2) v

1 * 2 = (1 + 2) v

V = 0.667 m/s

Directed to the east

========

Comment before rate in case any doubt, will reply for sure.. Goodluck